Thursday 09/13/12

9:00 A – 4:00P, + remotely 5:00P-7:20P, 8:20P-11:20PM

  • O/N STORM run still going smooth and in focus at 8:00 AM.
  • O/N STORM run completes on time, 9:30AM.  Break down run turn off scope.
  • Coding matlab wrapper for python / DaoSTORM (see post)
  • Rewrote insightM wrapper into smarter fxn based on new DaoSTORM wrapper.   Also has proper vargin inputs.
  • Initialize new local code repository Matlab_STORM for good copies of Matlab scripts for STORM analysis.
  • setup Matlab_STORM on Github
  • send git intro to Jeff.
  • screwed up pointer to DaoSTORM parameters file for test run 1.  Fixed and re-running analysis to compare.
  • Analysis completed.  Didn’t include tracking (repeat this?)
  • Attempting to get DaoSTORM running on Monet.  Need to recompile c libraries for Windows 7 Ultimate (I think this is the error.  Windows 2008 Server binaries not working).
  • Installing MinGW.
  • Downloaded windows compiled Lapack from here
  • Failed to install a functional version of 4 bit MinGW compiler (can download but  not install / compile to create executable form).  Attempts to change file in shell script and execute fail.
  • unable to link lapack.
  • Probably should be doing this with the 64 bit compiler anyway.
  • 5 hours of downloading installing and surfing for solutions.   No progress.  Yay windows for having STUPID DIFFICULT ability to get a simple 64 bit c-compiler working.   Try on Linux laptop next?
Fly work / planning
  • New fly stocks ordered
    • 2311 sna[18]
    • 33402 SuZ(12) TRiP valium20
    • 35389 Scm TRiP valium22
    • 36070 Pc 1 TRiP valium22
    • 33964 Pc 2 TRiP valium20
    • 33622 Pc 3 TRiP valium20
    • 35207 Ph-p TRiP valium22
    • 33669 Ph-p 2 TRiP valium20
    • 35297 Psc TRiP valium22
    • 35381 mBT TRiP valium22
    • 36786 Rad21 TRiP valium22
    • 25374 Act5-Gal4 w.
  • Check: previously ordered fly stocks have not arrived.
  • Sent introduction letter to Fred Bantignies, with demo figure.
  • Sent update on OligoPAINT STORM to Ting Wu.
snail project
Additional Notes
  • Colenso diagnosing issues with STORM4. Laser alignment on laser table substantially off-base, and beam alignement after fiber into objective substantially skewed, resulting in low laser power.  Should now be fixed.  Beam expansion lens now 100 mm focal point (smaller than previous).


This entry was posted in Summaries and tagged , . Bookmark the permalink.